OpenStudy (anonymous):

What is the product of 2√u(√5 +3√45)? simplify

OpenStudy (anonymous):

Can you rewrite it using the equation editor?

OpenStudy (anonymous):

whats that?

OpenStudy (anonymous):

nvm found it

OpenStudy (anonymous):

For instance, is it: \(2\sqrt{u}(\sqrt{5} + 3\sqrt{45})\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok. Let's first simplify what's inside the parentheses. \[2\sqrt{u}(\sqrt{5} + 3\sqrt{9 \times 5})\]\[2\sqrt{u}(\sqrt{5} + 3 \times 3\sqrt{5})\]Can you finish up here?

OpenStudy (anonymous):

i understand that part nut i get lost when i combine it with the variable. can you combine the √5 with the 2√u? i should be able to figure it out with a quick example if you don't mind.

OpenStudy (anonymous):

Alright. Before we get to that, can you simplify \(\sqrt{5} + s \times 3\sqrt{5}\)?

OpenStudy (anonymous):

i can't tell what the number being multiplied by 3 is. looks like an 8 maybe? in that it?

OpenStudy (anonymous):

*is

OpenStudy (anonymous):

Oh. That s should be a 3.

OpenStudy (anonymous):

9√10?

OpenStudy (anonymous):

and then what about the \(\sqrt{5}\)?

OpenStudy (anonymous):

im gonna go with 2√3 but 99% sure I'm wrong... i have like no idea how to do this radical stuff.

OpenStudy (anonymous):

\[\sqrt{5} + 9\sqrt{5} = 10\sqrt{5}\]

OpenStudy (anonymous):

So what we have at this point is: \[2\sqrt{u} \times 10\sqrt{5}\]

OpenStudy (anonymous):

What do you think?

OpenStudy (anonymous):

ok that makes sense combining like terms.

OpenStudy (anonymous):

Here's an example. \[5\sqrt{3x} \times 7\sqrt{2} \implies 35\sqrt{6x}\]

OpenStudy (anonymous):

so then it should be 20√u5???

OpenStudy (anonymous):

wow i was way off

OpenStudy (anonymous):

oh it was an example didn't see that part.

OpenStudy (anonymous):

Yeah. It would be \[20\sqrt{5u}\]Remember that the numbers always come before the variables :)

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

np :)